# Compositional Generative Multiphysics and Multi-component Simulation

### v1

@@ 3.1 MULTIPHYSICS SIMULATION
Our proposed framework of multiphysics simulation in Section 3.1 and multi-component simulation in Section 3.2, constitute our full method of compositional MultiPhysics and Multi-component Simulation with Diffusion models (MultiSimDiff). It circumvents the development of coupled programs that requires huge development efforts, and achieves multiphysics and multi-component simulation by composing the learned conditional energy functions according to the variable dependencies. Below, we test our method’s capability in challenging engineering problems.
@@ 4 EXPERIMENTS
In the experiments, we aim to answer the following questions: (1) Can MultiSimDiff predict coupled solutions (accounting for interactions between different physical processes) from models trained in decoupled data (assuming other processes are known and focus on solving a single process)? (2) Can MultiSimDiff predict large structure solutions from a model trained in small structure data? (3) Can MultiSimDiff outperform surrogate model [^2] The surrogate model mentioned in this article involves inputting the system’s input into a neural network to directly predict the output in one step, distinguishing it from the diffusion model, which uses the input as a condition to step-by-step generate the output through denoising. in both tasks? To answer these questions, we conduct experiments to assess our algorithm’s performance on two problems across three scenarios. In Section 4.1, we solve the reaction-diffusion equation. While it’s not a classic multiphysics coupling issue since both quantities are part of concentration fields, we consider them as separate physical processes to validate the capability of MultiSimDiff for multiphysics simulation. Section 4.2 examines a more complex scenario involving various types of coupling: region, interface, strong, weak, unidirectional, and bidirectional to further test the algorithm’s capacity to handle multiphysics simulation. Section 4.3 simplifies actual engineering problems to evaluate the algorithm’s effectiveness with multi-component simulation. Each experiment uses two network architectures, training both with their respective diffusion and surrogate models for comparison, employing consistent hyperparameters and settings to ensure fairness. The computational domains of experiment 1 and experiment 2 are on regular meshes, using Fourier neural operator (FNO) (Li et al., 2021; Lim et al., 2023) and U-Net (Ronneberger et al., 2015) as network architectures. The computational domain of experiment 3 is on a finite element mesh, using Geo-FNO (Li et al., 2023) and Transolver (Wu et al., 2024a) as the network architecture. Additionally, in experiment 3, we also compare MultiSimDiff with graph neural networks GIN (Xu et al., 2019) and Graph transformer SAN (Kreuzer et al., 2021). To highlight the difference between training and testing data, Appendix G calculates the Wasserstein distances between decoupled and coupled data in multiphysics problems, as well as between small and large structural data in multi-component problem, and visualizes them. The code is available at the repository. As another contribution to the community, we will also open-source the data to facilitate future method development of multiphysics and multi-component simulations. In Appendix K, we conduct a comparison of our dataset with existing scientific datasets (Takamoto et al., 2022).
@@ 4.1 Reaction-diffusion
Reaction-diffusion (RD) equations have found wide applications in the analysis of pattern formation, including chemical reactions. This experiment uses the 1D FitzHugh-Nagumo reaction-diffusion equation (Rao et al., 2023), it has two concentration fields: $u,v$. The objective is to predict the system’s evolution under different initial conditions. We use surrogate models that iteratively interact as a baseline. The training data consists of decoupled data, where other physical processes are assumed and treated as inputs to solve the equations governing the current physical process . In this experiment, a Gaussian random field (Bardeen et al., 1986) is employed to generate the other physical processes and initial conditions, and numerical algorithms are used to compute the solution of the current physical process. The validation data similarly consists of decoupled data not used during training. For the test data, initial conditions for both physical processes are generated using a Gaussian random field, and the ground-truth coupled solution is obtained using a fully coupled algorithm. Further details on the datasets, equation, network architecture, and training process are provided in Appendix B.
@@
MultiSimDiff (ours) + U-Net 0.0119 0.0141 0.0046 0.0174
Table 1 presents the relative L2 norm (L2 norm of prediction error divided by L2 norm of the ground-truth) in predictions made by surrogate model and MultiSimDiff on a validation set of decoupled data and a test set of coupled data. For FNO, the prediction error for $u$ is comparable between MultiSimDiff and surrogate models on decoupled data; however, MultiSimDiff shows a significantly larger error in predicting $v$, which is four times that of the surrogate model. As a result, the error in predicting the coupled solution for $v$ is greater than that of the surrogate model, while the error for $u$ is lower. For U-Net, the prediction errors for $u$ and $v$ are similar between MultiSimDiff and surrogate models on decoupled data, but MultiSimDiff achieves a lower error for the coupled solution.
This straightforward experiment tests the correctness of MultiSimDiff but shows no significant advantages over the surrogate model. However, the surrogate model fails in solving more complex problems, which will be discussed in the next section.
@@ 4.2 Nuclear thermal coupling
(a) Geometric structure.
(b) Coupling relationship between different physical processes.
Figure 2: Problem description of nuclear thermal coupling.
This experiment tests the performance of MultiSimDiff in more physical processes and coupling modes, including both regional and interface coupling, strong coupling and weak coupling, unidirectional and bidirectional coupling. We focus on nuclear thermal coupling in transient conditions for plate fuel elements. To simplify, a typical pin cell is analyzed (as shown in Fig. 2(a)), and the transient disturbance is modeled as a change in neutron flux density at the boundary. We aim to solve the neutron physics equation across the entire domain, the heat conduction equation in solid, and the flow heat transfer equations in fluid. This problem involves conjugate heat transfer between solid and fluid phases, presenting an interface coupling issue. Additionally, the negative feedback between the neutron physics field and the temperatures of the fluid and solid introduces a regional coupling. Aside from the unidirectional coupling from the fluid fields to the neutron physics field, all other interactions are bidirectional. The neutron physics field is weakly coupled with the temperature fields of solid and fluid, while the coupling effect at the interface between solid and fluid is strong. The objective is to predict the evolution of the entire physical system under different neutron boundary conditions. The coupling relationship between different physical processes are shown in Fig. 2(b).
@@
Figure 3: Comparison of prediction results between MultiSimDiff and surrogate model. The surrogate model fails on the test set of the coupled scenario.
Table 2 displays the relative prediction errors of surrogate models and MultiSimDiff on a validation set of decoupled data and a test set of coupled data. In single physical process prediction (decoupled data), surrogate models outperform MultiSimDiff. However, in predicting the coupled solution, all surrogate models fail except for the neutron physics field, with the predicted solid and fluid temperature fields shown in Fig. 3 (for more visualizations, see Fig. 5). The neutron physics field remains relatively accurate because the feedback from solid and fluid temperatures is weak and primarily driven by external input boundary conditions. In contrast, solid temperature and fluid fields are significantly influenced by other physical processes , leading to non-physical predictions due to the lack of iterative process data during training. In comparison, MultiSimDiff more accurately captures the morphology of coupled solutions and demonstrates higher accuracy. In addition, we further use DDIM (Song et al., 2021) to accelerate sampling and compare the operational efficiency of different methods. MultiSimDiff achieves an acceleration of up to 29 times, with detailed information in Appendices H and I.
@@
Figure 4: Problem setup of the prismatic fuel element. In (a), the left figure shows the entire reactor, with the purple section representing the reactor core, which mainly contains fuel elements, a matrix, and heat pipes. The right figure illustrates a portion of the reactor core, displaying only the matrix while omitting the heat pipes and fuel elements. This structure composed of 16 fuel elements is used to generate training data. (b) is a large structure composed of 64 elements used for testing.
This experiment tests the ability of MultiSimDiff to solve multi-component simulation problems, focusing on the thermal and mechanical analysis of prismatic fuel elements for a new type of reactor (Ma et al., 2022), as shown in Fig. 4. The reactor core consists of three components: fuel, matrix, and heat pipe. Since engineering focuses mainly on the matrix, we consider the fuel and fluid as boundary conditions for analysis. Different heat fluxes will be assigned to the fuel boundary to simulate various heat release behaviors of the fuel rods. The aim is to train a model that predicts its temperature $T$ and strain $\varepsilon_{x},\varepsilon_{y}$ based on the solutions of its three neighbors and its heat flux, and then use this basic model to predict larger structure as shown in Fig. 4(b).
The training data originate from a medium structure simulation that includes 16 fuel elements, as shown in Fig. 4(a); hence, a single simulation data point can generate 16 training data. The well-trained model will be tested on two structures: one is the medium structure used for data generation, and the other is a large structure containing 64 fuel elements. Further details on the datasets, network architecture, and training process are provided in Appendix D.
Table 3: Relative L2 norm of prediction error on prismatic fuel element experiment, for single-component and multi-component simulation. The multi-component includes 16-component (medium) and 64-component (large) simulations. The unit is $1\times 10^{-2}$.
@@
MultiSimDiff (ours) + Transolver 0.107 0.303 0.213 1.03 0.759 1.94
Table 3 presents the prediction relative errors of surrogate model and MultiSimDiff across three tasks: a single fuel element, a medium structure of 16 fuel elements, and a large structure of 64 fuel elements. The average relative error of strain $\varepsilon_{x},\varepsilon_{y}$ is denoted as $\varepsilon$. GIN and SAN learn on small graphs with 16 components and test on large graphs with 64 components. Due to the uniformity of graph structures in all training data and the fact that SAN learns a global relationship, SAN fails to predict larger structures. In contrast, GIN, capable of learning a local relationship, succeeds in handling larger structures. However, when compared to the surrogate model and MultiSimDiff, GIN has a larger error.
Subsequently, a comparative analysis between the surrogate model and MultiSimDiff has been conducted. The surrogate model performs better in predicting a single component, but for medium structure, MultiSimDiff outperforms it. It’s important to note that the surrogate model’s predictions occasionally diverge, necessitating adjustments to the relaxation factor to maintain stability. For the large structure, U-Net in the surrogate model demonstrates better stability, while the FNO model continues to diverge even after relaxation factor adjustments. MultiSimDiff is very stable and accurate, and no divergence phenomenon has been observed. Compared with the surrogate model, the relative error of MultiSimDiff has been reduced by 40.3% on average, demonstrating its ability to generalize to much larger multi-component simulations while trained on single components. In addition, we further use DDIM to accelerate sampling and compare the operational efficiency of different methods.MultiSimDiff achieves an acceleration of up to 41 times, with detailed information in Appendices H and I.
@@ 5 Limitation and future work
There are also several limitations of our proposed MultiSimDiff that provide exciting opportunities for future work. Firstly, in multiphysics simulation, although the MultiSimDiff trained on decoupled data can predict coupled solutions more accurately than baseline surrogate models, the prediction errors are still higher compared to single physical processes predictions. In addition, there is a certain gap in accuracy compared to models trained through coupled data, as shown in Appendix E. Future efforts can focus on improving dataset generation, training methods, and incorporating physical information to boost accuracy. Secondly, we plan to explore additional accelerated sampling algorithms, aiming to significantly improve efficiency while maintaining prediction accuracy. Lastly, the experiments in this paper simplify many aspects compared to real engineering problems, and future work will aim to validate the algorithm in more complex real-world scenarios.
@@ 6 CONCLUSION
This work presents MultiSimDiff as a novel method for multiphysics and multi-component simulations. In multiphysics scenarios, models trained on decoupled data can predict coupled solutions, while in multi-component simulations, models trained on small structures can extrapolate to larger ones. We develop three datasets to validate MultiSimDiff and compare it to the surrogate model method. Results show that MultiSimDiff effectively predicts coupled solutions in multiphysics simulations where surrogate models fail, and exhibits greater accuracy in predicting larger structures in multi-component simulations. We believe this approach provides a new approach to address multiphysics and multi-component simulations, important across science and engineering.
@@
Algorithm 2 Algorithm for multi-component simulation by MultiSimDiff
@@
3:     $\hat{z}_{v_{i}}^{e}\leftarrow z_{v_{i}}^{e}$ // update previous estimated solutions for each component $\hat{z}_{v_{i}}^{e}$
4:     $z_{v_{i}}^{e}\sim\mathcal{N}(0,\bf I)$ // initialize current estimated solutions for each component $z_{i}^{e}$
5:     $z_{{v_{i}},S}\sim\mathcal{N}(0,\bf I)$ // initialize solutions for each component $z_{v_{i}}$// denoising cycle of diffusion model:
6:     for $s=S,...,1$ do
7:        $\lambda=1-\frac{s}{S}$ if $k>1$ else 1 // define the weights of $\hat{z}_{i}^{e}$ and $z_{i}^{e}$
8:        $z_{\partial v}=f(z_{v_{1}}^{e},...,z_{v_{n}}^{e},\mathrm{adj})$ // update the solutions of surrounding components for each $z_{v_{i}}$
9:        $w\sim\mathcal{N}(0,\bf I)$// use weighted estimated solutions as conditions for single step denoising, // update all components together:
10:        $z_{v_{i},s-1}=\frac{1}{\sqrt{\alpha_{s}}}(z_{v_{i},s}-\frac{1-\alpha_{s}}{\sqrt{1-\overline{\alpha}_{s}}}\epsilon_{\theta}(z_{v_{i},s}\mid\lambda z_{\partial v_{i}}^{e}+(1-\lambda)\hat{z}_{\partial v_{i}}^{e},C))+\sigma_{s}w$// update the estimated solutions of all components together
11:        $z_{v_{i}}^{e}=\frac{1}{\sqrt{\overline{\alpha}_{s}}}(z_{{v_{i}},s}-\sqrt{1-\overline{\alpha}_{s}}\epsilon_{\theta}(z_{v_{i},s}\mid\lambda z_{v_{i}}^{e}+(1-\lambda)\hat{z}_{\partial v_{i}}^{e},C))$
12:     end for

### v2

@@ 3.2 Multi-component simulation
Our proposed framework of multiphysics simulation in Section 3.1 and multi-component simulation in Section 3.2, constitute our full method of compositional Multiphysics and Multi-component PDE Simulation with Diffusion models (M2PDE). It circumvents the development of coupled programs that requires huge development efforts, and achieves multiphysics and multi-component PDE simulation by composing the learned conditional energy functions according to the variable dependencies. Below, we test our method’s capability in challenging engineering problems.
@@ 4 Experiments
In the experiments, we aim to answer the following questions: (1) Can M2PDE predict coupled solutions (accounting for interactions between different physical processes) from models trained in decoupled data (assuming other processes are known and focus on solving a single process)? (2) Can M2PDE predict large structure solutions from a model trained in small structure data? (3) Can M2PDE outperform surrogate model [^2] The surrogate model in this article refers to a neural network that directly maps system inputs to outputs in one step, unlike diffusion models that generate outputs step-by-step through denoising. For multiphysics problems, surrogate models are trained separately for each physical field using decoupled data, and obtain coupled solutions through iteration.. For multi-component problems, surrogate models use domain decomposition concepts, predicting each component iteratively based on surrounding components to achieve convergence. in both tasks? To answer these questions, we conduct experiments to assess our algorithm’s performance on two problems across three scenarios. In Section 4.1, we solve the reaction-diffusion equation. While it’s not a classic multiphysics coupling issue since both quantities are part of concentration fields, we consider them as separate physical processes to validate the capability of M2PDE for multiphysics simulation. Section 4.2 explores a complex nuclear engineering scenario, involving various types of coupling: region, interface, strong, weak, unidirectional, and bidirectional to further test the algorithm’s capacity to handle multiphysics simulation. Section 4.3 simplifies actual nuclear engineering problems to evaluate the algorithm’s effectiveness with multi-component simulation. Each experiment uses two network architectures, training both with their respective diffusion and surrogate models for comparison, employing consistent hyperparameters and settings to ensure fairness. The computational domains of experiment 1 and experiment 2 are on regular meshes, using Fourier neural operator (FNO) (Li et al., 2021; Lim et al., 2023) and U-Net (Ronneberger et al., 2015) as network architectures. The computational domain of experiment 3 is on a finite element mesh, using Geo-FNO (Li et al., 2023) and Transolver (Wu et al., 2024a) as the network architecture. Additionally, in experiment 3, we also compare M2PDE with graph neural networks GIN (Xu et al., 2019) Graph transformer SAN (Kreuzer et al., 2021) and MeshGraphNet (Pfaff et al., 2020).
To highlight the difference between training and testing data, Appendix G calculates the Wasserstein distances between decoupled and coupled data in multiphysics problems, as well as between small and large structural data in multi-component problem, and visualizes them. In Appendix K, we conduct a comparison of our dataset with existing scientific datasets (Takamoto et al., 2022). As another contribution to the community, we will also open-source the data to facilitate future method development of multiphysics and multi-component PDE simulations.
@@ 4.1 Reaction-diffusion
Reaction-diffusion (RD) equations have found wide applications in the analysis of pattern formation, including chemical reactions. This experiment uses the 1D FitzHugh-Nagumo reaction-diffusion equation (Rao et al., 2023), it has two concentration fields: $u,v$. The objective is to predict the system’s evolution under different initial conditions. The training data consists of decoupled data, where other physical processes are assumed and treated as inputs to solve the equations governing the current physical process . Gaussian random field (Bardeen et al., 1986) is employed to generate the other physical processes and initial conditions, and numerical algorithms are used to compute the solution of the current physical process. The validation data similarly consists of decoupled data not used during training. The test data consists of coupled solutions obtained using fully coupled algorithms. Further details on the datasets, equation, network architecture, and training process are provided in Appendix B.
@@
M2PDE (ours) + U-Net 0.0119 0.0141 0.0046 0.0174
Table 1 presents the relative L2 norm (L2 norm of prediction error divided by L2 norm of the ground-truth) in predictions made by surrogate model and M2PDE on a validation set of decoupled data and a test set of coupled data. For FNO, the prediction error for $u$ is comparable between M2PDE and surrogate models on decoupled data; however, M2PDE shows a significantly larger error in predicting $v$, which is four times that of the surrogate model. As a result, the error in predicting the coupled solution for $v$ is greater than that of the surrogate model, while the error for $u$ is lower. For U-Net, the prediction errors for $u$ and $v$ are similar between M2PDE and surrogate models on decoupled data, but M2PDE achieves a lower error for the coupled solution.
This straightforward experiment tests the correctness of M2PDE but shows no significant advantages over the surrogate model. However, the surrogate model fails in solving more complex problems, which will be discussed in the next section.
@@ 4.2 Nuclear thermal coupling
Figure 2: Geometric structure (left); Coupling relationship between different physical processes (right).
This experiment evaluates M2PDE’s performance in handling various physical processes and coupling modes, including regional and interface coupling, strong and weak coupling, and unidirectional and bidirectional coupling. The focus is on nuclear-thermal coupling in transient conditions for plate fuel elements, using a simplified pin cell analysis with a transient disturbance modeled as a change in neutron flux density at the boundary. This physical system requires solving neutron physics, heat conduction, and flow heat transfer equations, taking into account negative feedback between neutron physics and temperature, unidirectional coupling from fluid to neutron field, and strong interface coupling between solid and fluid. The goal is to predict the system’s evolution under different neutron boundary conditions. The geometric and coupling relationships are illustrated in Fig. 2.
@@
Figure 3: Comparison of prediction results between M2PDE and surrogate model. The surrogate model fails on the test set of the coupled scenario.
Table 2 displays the relative prediction errors of surrogate models and M2PDE on a validation set of decoupled data and a test set of coupled data. In single physical process prediction (decoupled data), surrogate models outperform M2PDE. However, in predicting the coupled solution, all surrogate models fail except for the neutron physics field, with the predicted solid and fluid temperature fields shown in Fig. 3 (for more visualizations, see Fig. 6). The neutron physics field remains relatively accurate because the feedback from solid and fluid temperatures is weak and primarily driven by external input boundary conditions. In contrast, solid temperature and fluid fields are significantly influenced by other physical processes, leading to non-physical predictions due to the lack of iterative process data during training. In comparison, M2PDE more accurately captures the morphology of coupled solutions and demonstrates higher accuracy. In addition, we further use DDIM (Song et al., 2021) to accelerate sampling and compare the operational efficiency of different methods. M2PDE achieves an acceleration of up to 29-fold speedup, with detailed information in Appendices H and I.
@@ 4.3 Prismatic fuel element
This experiment tests the ability of M2PDE to solve multi-component simulation problems, focusing on the thermal and mechanical analysis of prismatic fuel elements for a new type of reactor (Ma et al., 2022). The reactor core consists of three components: fuel, matrix, and heat pipe. Since engineering focuses mainly on the matrix, we consider the fuel and fluid as boundary conditions for analysis. Different heat fluxes will be assigned to the fuel boundary to simulate various heat release behaviors of the fuel rods. The aim is to train a model that predicts its temperature $T$ and strain $\varepsilon_{x},\varepsilon_{y}$ based on the solutions of its three neighbors and its heat flux.
The training data originate from a medium structure simulation that includes 16 fuel elements; hence, a single simulation data point can generate 16 training data. The well-trained model will be tested on two structures: one is the medium structure used for data generation, and the other is a large structure containing 64 fuel elements, as shown in Fig. 4. Further details on the datasets, network architecture, and training process are provided in Appendix D.
@@
Figure 4: In (a), the left figure shows the entire reactor. The right figure illustrates a portion of the reactor core. This structure composed of 16 fuel elements is used to generate training data. (b) is a large structure composed of 64 elements used for testing.
Table 3: Relative L2 norm of prediction error on prismatic fuel element experiment, for single-component and 16-component (medium) and 64-component (large) simulations. The unit is $1\times 10^{-2}$.
@@
M2PDE (ours) + Transolver 0.144 0.339 0.288 1.01 0.793 2.07
Table 3 presents the prediction relative errors of surrogate model and M2PDE across three tasks: a single fuel element, a medium structure of 16 fuel elements, and a large structure of 64 fuel elements. The average relative error of strain $\varepsilon_{x},\varepsilon_{y}$ is denoted as $\varepsilon$. GIN, SAN, and MeshGraphNet learn on small graphs with 16 components and test on large graphs with 64 components. Due to the uniformity of graph structures in all training data and the fact that SAN learns a global relationship, SAN fails to predict larger structures. GIN and MeshGraphnet are capable of learning local relationships, thereby successfully handling larger-scale structures. Specifically, MeshGraphnet achieves the lowest error when learning structures with 16 components. However, when extrapolating to even larger structures, M2PDE exhibits lower error than MeshGraphnet.
Subsequently, a comparative analysis between the surrogate model and M2PDE has been conducted. The surrogate model performs better in predicting a single component, but for medium structure, M2PDE outperforms it. It’s important to note that the surrogate model’s predictions occasionally diverge, necessitating adjustments to the relaxation factor to maintain stability. For the large structure, U-Net in the surrogate model demonstrates better stability, while the FNO model continues to diverge even after relaxation factor adjustments. M2PDE is very stable and accurate, and no divergence phenomenon has been observed. In term of strain, the relative error of M2PDE has been reduced by 41% Compared with the surrogate model. In addition, we further use DDIM to accelerate sampling and compare the operational efficiency of different methods. M2PDE achieves an acceleration of up to 41-fold speedup, with detailed information in Appendices H and I.
@@ 5 Limitation and future work
There are also several limitations of our proposed M2PDE that provide exciting opportunities for future work. Firstly, in multiphysics simulation, although the M2PDE trained on decoupled data can predict coupled solutions more accurately than baseline surrogate models, the prediction errors are still higher compared to single physical processes predictions. In addition, there is a certain gap in accuracy compared to models trained through coupled data, as shown in Appendix E. Future efforts can focus on improving dataset generation, training methods, and incorporating physical information to boost accuracy. Secondly, we plan to explore additional accelerated sampling algorithms, aiming to significantly improve efficiency while maintaining prediction accuracy. Lastly, while the experiments in this paper have simplified certain aspects of real-world engineering problems to focus on key issues and improve research efficiency, future work will aim to apply this method to more complex practical engineering problems to further validate its effectiveness and practicality.
@@ 6 Conclusion
This work presents M2PDE as a novel method for multiphysics and multi-component PDE simulations, driven by the needs of real-world engineering applications. In multiphysics scenarios, models trained on decoupled data can predict coupled solutions, while in multi-component simulations, models trained on small structures can extrapolate to larger ones. We develop three datasets to validate M2PDE and compare it to the surrogate model method. Results show that M2PDE effectively predicts coupled solutions in multiphysics simulations where surrogate models fail, and exhibits greater accuracy in predicting larger structures in multi-component simulations. We believe this approach provides a new approach to address multiphysics and multi-component PDE simulations, important across science and engineering.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@
Algorithm 2 Algorithm for multi-component simulation by M2PDE
@@
3:      // initialize each component $z_{v_{i}}$, for $i$ in $1,...,N$
4:     $\hat{z}_{v_{i}}^{e}\leftarrow z_{v_{i}}^{e}$ // update previous estimated solutions for each component $\hat{z}_{v_{i}}^{e}$
5:     $z_{v_{i}}^{e}\sim\mathcal{N}(0,\bf I)$ // initialize current estimated solutions for each component $z_{i}^{e}$
6:     $z_{{v_{i}},S}\sim\mathcal{N}(0,\bf I)$ // initialize solutions for each component $z_{v_{i}}$// denoising cycle of diffusion model:
7:     for $s=S,...,1$ do
8:        $\lambda=1-\frac{s}{S}$ if $k>1$ else 1 // define the weights of $\hat{z}_{i}^{e}$ and $z_{i}^{e}$
9:        $z_{\partial v}=f(z_{v_{1}}^{e},...,z_{v_{N}}^{e},\mathrm{adj})$ // update the solutions of surrounding components for each $z_{v_{i}}$
10:        $w\sim\mathcal{N}(0,\bf I)$// use weighted estimated solutions as conditions for single step denoising, // update all components together:
11:        $z_{v_{i},s-1}=\frac{1}{\sqrt{\alpha_{s}}}(z_{v_{i},s}-\frac{1-\alpha_{s}}{\sqrt{1-\overline{\alpha}_{s}}}\epsilon_{\theta}(z_{v_{i},s}\mid\lambda z_{\partial v_{i}}^{e}+(1-\lambda)\hat{z}_{\partial v_{i}}^{e},C))+\sigma_{s}w$// update the estimated solutions of all components together
12:        $z_{v_{i}}^{e}=\frac{1}{\sqrt{\overline{\alpha}_{s}}}(z_{{v_{i}},s}-\sqrt{1-\overline{\alpha}_{s}}\epsilon_{\theta}(z_{v_{i},s}\mid\lambda z_{v_{i}}^{e}+(1-\lambda)\hat{z}_{\partial v_{i}}^{e},C))$